OpenStudy (anonymous):

A geometry student wants to draw a rectangle inscribed in the ellipse x^2 +4y^2 =9. What is the area of the largest rectangle that the student can draw?

OpenStudy (anonymous):

Let the lengths of the sides of rectangle be a*2 and b*2 parallel to x axis and y axis respectively; a^2 + 4b^2 = 9 b = (9- a^2)/4 Let the area of the rectangle be a function of a: A(a) = a*2*(9- a^2)/4*2 = 9a - a^3 you will want to find the local extrema of A(a), that is, dA/da = 0 9 - 3a^2 = 0 a = sqrt(3) or -sqrt(3)

OpenStudy (anonymous):

a second order derivative test shows that A''(sqrt(3)) < 0, A''(-sqrt(3)) > 0, therefore A(sqrt(3)) is a local maxima. The area of the rectangle is therefore 9*sqrt(3) - (sqrt(3))^3

OpenStudy (anonymous):

Why are the side lengths equal to \[a ^{2} and. b ^{2}\]?

OpenStudy (anonymous):

not a^2 , a*2 

OpenStudy (anonymous):

OH. That makes a lot more sense XD Thanks!

OpenStudy (ranga):

If a^2 + 4b^2 = 9 Shouldn't b = sqrt( (9- a^2)/4 ) ?

OpenStudy (ranga):

I am getting: \[a = \frac{ 3\sqrt{2} }{ 2 }\] \[b = \frac{ 3\sqrt{2} }{ 4 }\]

OpenStudy (anonymous):

b = sqrt((9- a^2)/4) Let the area of the rectangle be a function of a: A(a) = a*2*((9- a^2)/4)^(1/2)*2 = a*2*((9- a^2)*(1/2)) you will want to find the local extrema of A(a), that is, dA/da = 0 2*(9-a^2)^(1/2)-2*a^2*(9-a^2)^(-1/2)=0 a = 3/sqrt(2) b = 9/8

OpenStudy (ranga):

I am getting Maximum Area = 9 Area = 4ab with a and b given in my earlier reply.

OpenStudy (anonymous):

DX the last line of my solution was supposed to be b^2 I should really be more careful with these trivial things.

OpenStudy (ranga):

yeah that is b^2. Your a and mine are the same if you multiply your numerator and denominator by sqrt(2) And if you take the square root of 9/8 your b will be the same as mine. Some textbooks and teachers don't like radical signs in the denominator and so I get rid of them in the denominator.